# Super-Linear: A Lightweight Pretrained Mixture of Linear Experts for Time Series Forecasting

### v1

@@ Appendix B Theoretical Analysis
Appendix C Datasets
C.1 Pre-training Datasets
@@ C.2 Evaluation Datasets
is a benchmark designed to evaluate the generalization ability of time series forecasting models across highly diverse and heterogeneous datasets. Unlike domain-specific benchmarks, it aggregates time series from a wide range of sources—including energy, finance, healthcare, traffic, and climate—covering different sampling rates, lengths, and dynamics. The goal is to test whether models trained in a universal or foundation-style setting can transfer effectively to unseen distributions without task-specific fine-tuning. By emphasizing cross-domain robustness, scalability, and zero-shot forecasting performance, GIFT-Eval serves as a more realistic and challenging measure of a model’s ability to act as a general-purpose time series forecaster. A complete description of all the datasets is given in 11.
@@
BizITObs - App (Palaskar et al., 2024) Web/CloudOps 10sec 1 8,834 8,834 8,834 8,834 2 60 15 600 2 900 1
BizITObs - Service (Palaskar et al., 2024) Web/CloudOps 10sec 21 8,835 8,835 8,835 185,535 2 60 15 600 2 900 1
BizITObs - L2C (Palaskar et al., 2024) Web/CloudOps 5min 1 31,968 31,968 31,968 31,968 7 48 20 480 7 720 5
BizITObs - L2C (Palaskar et al., 2024) Web/CloudOps hourly 1 2,664 2,664 2,664 2,664 7 48 6 480 1 720 1
Bitbrains - Fast (Shen et al., 2015) Web/CloudOps 5min 1,250 8,640 8,640 8,640 10,800,000 2 48 18 480 2 720 2
Bitbrains - Fast (Shen et al., 2015) Web/CloudOps hourly 1,250 721 721 721 901,250 2 48 2 - - - -
Bitbrains - rnd (Shen et al., 2015) Web/CloudOps 5min 500 8,640 8,640 8,640 4,320,000 2 48 18 480 2 720 2
Bitbrains - rnd (Shen et al., 2015) Web/CloudOps hourly 500 720 720 720 360,000 2 48 2 - - - -
@@
M4 Yearly (Godahewa et al., 2021) Econ/Fin yearly-DEC 22,974 37 19 284 845,109 1 6 1 - - - -
M4 Quarterly (Godahewa et al., 2021) Econ/Fin quarterly-DEC 24,000 100 24 874 2,406,108 1 8 1 - - - -
M4 Monthly (Godahewa et al., 2021) Econ/Fin monthly 48,000 234 60 2,812 11,246,411 1 18 1 - - - -
M4 Weekly (Godahewa et al., 2021) Econ/Fin weekly-SUN 359 1,035 93 2,610 371,579 1 13 1 - - - -
M4 Daily (Godahewa et al., 2021) Econ/Fin daily 4,227 2,371 107 9,933 10,023,836 1 14 1 - - - -
M4 Hourly (Godahewa et al., 2021) Econ/Fin hourly 414 902 748 1,008 373,372 1 48 2 - - - -
Hospital (Godahewa et al., 2021) Healthcare monthly 767 84 84 84 64,428 1 12 1 - - - -
COVID Deaths (Godahewa et al., 2021) Healthcare daily 266 212 212 212 56,392 1 30 1 - - - -
US Births (Godahewa et al., 2021) Healthcare daily 1 7,305 7,305 7,305 7,305 1 30 20 - - - -
US Births (Godahewa et al., 2021) Healthcare weekly-TUE 1 1,043 1,043 1,043 1,043 1 8 14 - - - -
US Births (Godahewa et al., 2021) Healthcare monthly 1 240 240 240 240 1 12 2 - - - -
Saugeen (Godahewa et al., 2021) Nature daily 1 23,741 23,741 23,741 23,741 1 30 20 - - - -
Saugeen (Godahewa et al., 2021) Nature weekly-THU 1 3,391 3,391 3,391 3,391 1 8 20 - - - -
Saugeen (Godahewa et al., 2021) Nature monthly 1 780 780 780 780 1 12 7 - - - -
Temp. Rain (Godahewa et al., 2021) Nature daily 32,072 725 725 725 780 1 30 3 - - - -
KDD Cup 2018 (Godahewa et al., 2021) Nature hourly 270 10,898 9,504 10,920 2,942,364 1 48 20 480 2 720 2
KDD Cup 2018 (Godahewa et al., 2021) Nature daily 270 455 396 455 122,791 1 30 2 - - - -
Car Parts (Godahewa et al., 2021) Sales monthly 2,674 51 51 51 136,374 1 12 1 - - - -
@@
C.3 Dataset Training Details
@@
During training, we uniformly sample from the pool of target frequencies (see Appendix C.3), independently of the dataset source. This reflects our focus on frequency diversity rather than dataset diversity. For each dataset, we use the first 80% of its timeline for training and reserve the remaining 20% for validation. Since datasets vary in channel dimensionality and length, we adopt channel-independent (i.e., univariate) training, treating each channel as a separate time series.
Appendix D Benchmark Models
D.1 Pretrained Models
@@
D.2 In-Domain Models
@@
Appendix E Super-Linear Frequency Expert Weights
@@
Appendix F Hyperparameters
In this section we present the hyper parameter for the pretrained ZS Super-Linear 12 and in-domain (FS) 13.
@@ Appendix F Hyperparameters
Appendix G Limitations
Despite its efficiency and strong performance, Super-Linear has several limitations. First, its reliance on linear representations may restrict its ability to capture complex temporal dependencies or nonlinear patterns that are common in many real-world time series. While the frequency-specialized design helps mitigate this to some extent, it may still fall short in scenarios where nonlinear interactions between frequencies are crucial. Second, although Super-Linear can process relatively long input sequences due to its lightweight architecture, its context modeling is limited by 512 lookback steps, which may hinder performance on tasks that require reasoning over very long lookback contexts. Finally, the model may struggle with time series that exhibit uncommon or transient frequency components not well-aligned with the predefined frequencies in 12. Expert specialization relies on fixed frequency filters, which may struggle to represent rare or shifting frequency patterns, potentially hindering generalization in such scenarios. While complementary layers aim to bridge information gaps—particularly for uncommon frequencies—this limitation may still persist.

### v2

@@ Appendix C Theoretical Analysis
Appendix D Datasets
D.1 Pre-training Datasets
@@ D.2 Evaluation Datasets
is a benchmark designed to evaluate the generalization ability of time series forecasting models across highly diverse and heterogeneous datasets. Unlike domain-specific benchmarks, it aggregates time series from a wide range of sources—including energy, finance, healthcare, traffic, and climate—covering different sampling rates, lengths, and dynamics. The goal is to test whether models trained in a universal or foundation-style setting can transfer effectively to unseen distributions without task-specific fine-tuning. By emphasizing cross-domain robustness, scalability, and zero-shot forecasting performance, GIFT-Eval serves as a more realistic and challenging measure of a model’s ability to act as a general-purpose time series forecaster. A complete description of all the datasets is given in 13.
@@
BizITObs - App Palaskar et al. (2024) Web/CloudOps 10sec 1 8,834 8,834 8,834 8,834 2 60 15 600 2 900 1
BizITObs - Service Palaskar et al. (2024) Web/CloudOps 10sec 21 8,835 8,835 8,835 185,535 2 60 15 600 2 900 1
BizITObs - L2C Palaskar et al. (2024) Web/CloudOps 5min 1 31,968 31,968 31,968 31,968 7 48 20 480 7 720 5
BizITObs - L2C Palaskar et al. (2024) Web/CloudOps hourly 1 2,664 2,664 2,664 2,664 7 48 6 480 1 720 1
Bitbrains - Fast Shen et al. (2015) Web/CloudOps 5min 1,250 8,640 8,640 8,640 10,800,000 2 48 18 480 2 720 2
Bitbrains - Fast Shen et al. (2015) Web/CloudOps hourly 1,250 721 721 721 901,250 2 48 2 - - - -
Bitbrains - rnd Shen et al. (2015) Web/CloudOps 5min 500 8,640 8,640 8,640 4,320,000 2 48 18 480 2 720 2
Bitbrains - rnd Shen et al. (2015) Web/CloudOps hourly 500 720 720 720 360,000 2 48 2 - - - -
@@
M4 Yearly Godahewa et al. (2021) Econ/Fin yearly-DEC 22,974 37 19 284 845,109 1 6 1 - - - -
M4 Quarterly Godahewa et al. (2021) Econ/Fin quarterly-DEC 24,000 100 24 874 2,406,108 1 8 1 - - - -
M4 Monthly Godahewa et al. (2021) Econ/Fin monthly 48,000 234 60 2,812 11,246,411 1 18 1 - - - -
M4 Weekly Godahewa et al. (2021) Econ/Fin weekly-SUN 359 1,035 93 2,610 371,579 1 13 1 - - - -
M4 Daily Godahewa et al. (2021) Econ/Fin daily 4,227 2,371 107 9,933 10,023,836 1 14 1 - - - -
M4 Hourly Godahewa et al. (2021) Econ/Fin hourly 414 902 748 1,008 373,372 1 48 2 - - - -
Hospital Godahewa et al. (2021) Healthcare monthly 767 84 84 84 64,428 1 12 1 - - - -
COVID Deaths Godahewa et al. (2021) Healthcare daily 266 212 212 212 56,392 1 30 1 - - - -
US Births Godahewa et al. (2021) Healthcare daily 1 7,305 7,305 7,305 7,305 1 30 20 - - - -
US Births Godahewa et al. (2021) Healthcare weekly-TUE 1 1,043 1,043 1,043 1,043 1 8 14 - - - -
US Births Godahewa et al. (2021) Healthcare monthly 1 240 240 240 240 1 12 2 - - - -
Saugeen Godahewa et al. (2021) Nature daily 1 23,741 23,741 23,741 23,741 1 30 20 - - - -
Saugeen Godahewa et al. (2021) Nature weekly-THU 1 3,391 3,391 3,391 3,391 1 8 20 - - - -
Saugeen Godahewa et al. (2021) Nature monthly 1 780 780 780 780 1 12 7 - - - -
Temp. Rain Godahewa et al. (2021) Nature daily 32,072 725 725 725 780 1 30 3 - - - -
KDD Cup 2018 Godahewa et al. (2021) Nature hourly 270 10,898 9,504 10,920 2,942,364 1 48 20 480 2 720 2
KDD Cup 2018 Godahewa et al. (2021) Nature daily 270 455 396 455 122,791 1 30 2 - - - -
Car Parts Godahewa et al. (2021) Sales monthly 2,674 51 51 51 136,374 1 12 1 - - - -
@@
D.3 Dataset Training Details
@@
The training complexity of Super-Linear, consists of two stages with runtimes $T_{1}$ and $T_{2}$ (see Fig. 4), for the stage 1 expert training and the stage 2 router + complementary experts training, respectively. Here, $T_{1}$ = $\frac{T_{expert}}{c}$, where $T_{\text{expert}}$ is the training time for a single expert and $c$ is the number of available GPUs.
On our system (RTX3090), we measure:
$T_{2}\approx$1.17hours, and $T_{expert}\approx$0.72hours
Data Resampling To ensure frequency diversity across the pre-trained datasets, we transform each dataset to span a wider range of frequencies, as illustrated in 11 and listed in 11. For each dataset, we sample a sequence and apply a scaling factor that either expands or compresses the signal, effectively shifting its dominant frequency. Since every dataset has a known dominant frequency, we can systematically rescale it to obtain sequences representing different frequencies.
For example, a dataset sampled at an hourly or daily rate with a dominant frequency of $\frac{1}{24}$ can be downsampled by a factor of $\frac{7}{24}$, yielding a sequence of length $\frac{7}{24}$ of the original. The same procedure applies for upsampling, where signals are expanded using a scaling factor greater than $1$. The maximal scaling factor used is $20$. We use linear interpolation as the resampling operator.
@@
During training, we uniformly sample from the pool of target frequencies (see Appendix D.3), independently of the dataset source. This reflects our focus on frequency diversity rather than dataset diversity. For each dataset, we use the first 80% of its timeline for training and reserve the remaining 20% for validation. Since datasets vary in channel dimensionality and length, we adopt channel-independent (i.e., univariate) training, treating each channel as a separate time series.
Appendix E Benchmark Models
E.1 Pretrained Models
@@
TiRex (Auer et al., 2025) TiRex is a zero-shot forecasting model built on xLSTM, a modern LSTM variant that combines strong state-tracking with competitive in-context learning. Its design is enhanced by Contiguous Patch Masking, which stabilizes long-horizon autoregressive predictions, and by tailored data augmentations that improve robustness. This architecture enables TiRex to outperform much larger transformer-based models on standard benchmarks.
E.2 In-Domain Models
@@
Appendix F Super-Linear Frequency Expert Weights
@@
Appendix G Hyperparameters
In this section we present the hyper parameter for the pretrained ZS Super-Linear 14 and in-domain (FS) 15.
@@ Appendix G Hyperparameters
Appendix H Limitations
Despite its efficiency and strong performance, Super-Linear has several limitations. First, its reliance on linear representations may restrict its ability to capture complex temporal dependencies or nonlinear patterns that are common in many real-world time series. While the frequency-specialized design helps mitigate this to some extent, it may still fall short in scenarios where nonlinear interactions between frequencies are crucial. Second, although Super-Linear can process relatively long input sequences due to its lightweight architecture, its context modeling is limited by 512 lookback steps, which may hinder performance on tasks that require reasoning over very long lookback contexts. Finally, the model may struggle with time series that exhibit uncommon or transient frequency components not well-aligned with the predefined frequencies in 11. Expert specialization relies on fixed frequency filters, which may struggle to represent rare or shifting frequency patterns, potentially hindering generalization in such scenarios. While complementary layers aim to bridge information gaps—particularly for uncommon frequencies—this limitation may still persist.